Profiles of Dialogue

Erik C.W. Krabbe

March 31, 1999

Abstract

According to Walton (1999: 53), a “profile of dialogue is a connected sequence of moves and countermoves in a conversational exchange of a type that is goal-directed and can be represented in a normative model of dialogue.” I should like to add that, generally, a profile is not to be identified with a single occurrence of a sequence, but is supposed to give us some general features of such sequences at some intermediate level of abstraction. Also a profile, generally, does not display just one (abstract) sequence, but a number of them, organized as branches in a tree. In this paper I hope to clarify the notion of “profile of dialogue” to a certain extent, by giving examples. There will be no attempt to reach a definition of the concept. At the end, I shall indicate how profiles can be useful for the study of argumentation.

Contents

1 Introduction 2

2 Why argue? 2

3 Why not? 3

4 Many Questions 4

5 As discussion starts 5

6 Profiles for equivocation criticism 7

7 Conclusion 11

1 Introduction

According to Walton (1999: 53), a “profile of dialogue is a connected sequence of moves and countermoves in a conversational exchange of a type that is goaldirected and can be represented in a normative model of dialogue.” I should like to add that, generally, a profile is not to be identified with a single occurrence of a sequence, but is supposed to give us some general features of such sequences at some intermediate level of abstraction. Also a profile, generally, does not display just one (abstract) sequence, but a number of them, organized as branches in a tree. In this paper I hope to clarify the notion of “profile of dialogue” to a certain extent, by giving examples. There will be no attempt to reach a definition of the concept. At the end, I shall indicate how profiles can be useful for the study of argumentation.

2 Why argue?

Presumably, argumentation is a goal-directed activity of some type. If so, one may ask what the goal is. Why do people argue? Actually, there would be no reason to argue if we all shared our opinions on all subjects. But it is a basic fact about the human condition that we do not. Even so, there would be no occasion for argument if people kept their opinions to themselves, but again they do not. And though sometimes one would prefer that silence had prevailed, it is in general a fortunate circumstance that people voice there opinions and disagreements, so that they can start arguing about them.

As soon as one party expresses a point of view vis-à-vis another party, and the other party expresses that this point of view is not as such accepted, a conflict is born (Barth & Krabbe 1982: 56, Def. 1). It must be stressed that the second party need not express an opposed point of view: it is sufficient, for there to be a conflict, if this party expresses nonacceptance (critical doubt). Critical doubt as to the acceptability of a statement is a position, but not a point of view (standpoint). Thus Peter may express the point of view that the fine skating weather is holding and be confronted by a mere critical “Why?” from the side of Olga. This type of conflict is called “pure” or “simple” and constitutes a basis for the analysis of all other types of conflict, for instance of cases where Peter is confronted by a contradictory or contrary statement. The pragma-dialecticians have further analyzed these types of conflict (or dispute) (Van Eemeren & Grootendorst 1992, Ch. 2).

Once a conflict, or dispute, has arisen, one may ask how to resolve it. Though there are many ways to settle a dispute (if a fight ensues, and one party dies, the dispute is no doubt settled; but one may also use peaceful means such as negotiation or casting a die), the only way to resolve a dispute is for each party to try to achieve a change in the position taken by the other through serious attempts to convince the other; i.e., through critical discussion (Van Eemeren & Grootendorst 1992: 34). The distinction between settling and resolving may be terminological, but is not, on that account, without significance. We need clear terms to briefly state our purpose in arguing.

3 Why not?

Now suppose that Peter and Olga expressed there positions as to the skating weather. And suppose they decided not to fight or negotiate, but to resolve their dispute by critical discussion. The following dialogue might ensue (this may also be called a concrete profile of dialogue):

- Peter: The fine skating weather is holding.

- Olga: Why?

- Peter: Why not?

How is one to judge Peter’s last move? Would it be acceptable? If no further context is given, the (more abstract) profile of dialogue, in this case, would be:

- P: p

- O: Why(p)?

- P: Why(not-p)?

It seems many things are wrong with this profile. First of all, there is a dialectical shift involved. Not a shift from one type of dialogue to another (say, negotiation), since the framework of critical discussion is not abandoned, but an internal shift within critical discussion. More precisely the shift (or attempted shift) is from a simple conflict in wich only P has a point of view to defend to a more complex type (more properly called “dispute”) in which O has to defend the opposite point of view. It’s an attempted shift of attitude (Walton & Krabbe 1995: 108). Even though dialectical shifts are not necessarily wrong or fallacious, this one seems to thwart the purpose of critical discussion. It seems that if P is allowed to reply “Why(not-p)?”, O must be granted the right to reply by “Why(p)?” again, and so on. Unless other rules prevent this, a resolution of the original conflict cannot be reached. Even if such an indefinitely protracted altercation is somehow ruled out , P’s move can still be characterized as “adynamic” in that it does not advance the process of resolution.

Moreover, in this profile there is a shifting of the burden of proof and a straw man fallacy. Hence there is an infringement of at least two pragmadialectical rules. Rule 2 says that a “party who advances a standpoint is obliged to defend it if the other party asks him to do so” (Van Eemeren & Grootendorst 1992: 117), and shifting the burden of proof is dicussed by Van Eemeren and Grootendorst as one of the ways to break this rule (120ff). Clearly, in the present profile, P is shifting the burden of proof to O. Rule 3 says that a “party’s attack on a standpoint must relate to the standpoint that has indeed been advanced by the other party” (125). Since no standpoint at all has been advanced by O, P’s attack “Why(not-p)?” breaks this rule as well. It is also possible to subsume the present case in the class of argumenta ad ignorantiam, since P seems to imply that, if O cannot prove her point, he has been shown to be in the right (Krabbe 1995: 256ff).

The why/why-not case is full of suggestions that may be profitably studied using both profiles of dialogue and pragma-dialectical rules. The more abstract profile puts the case on a kind of intermediate level of abstraction: though no reconstruction within a completely defined system of formal dialectic was undertaken, we were able to abstract from a number of specifics of the case and came to the conclusion that answering a “Why?” by a “Why not?” is not commendable as a contribution to critical discussion.

But then, the case might occur in a larger context so that P’s “Why(notp)?” move could be justifiable in some cases. Suppose, by way of example, that Peter and Olga agree about the weather, that it is common knowledge among them that the good skating weather is holding, so that this proposition can be used by either of them as a basic reason in argument. Suppose that, instead of discussing the weather, they now discuss when to go skating:

- Peter: We might postpone our trip until tomorrow.

- Olga: Why?

- Peter: The fine skating weather is holding.

- Olga: Why?

- Peter: Why not?

Here Peter may well ask why Olga suddenly retracts commitment to a piece of common knowledge. The abstract profile seems to look thus:

- P: p

- O: Why(p)?

- P: r

- O: Why(r)?

- P: Why(not-r)?

However, this profile still suggests that Peter is committing a straw man fallacy. O may have retracted commitment to the common presumption r, but there is no indication that she is committed to not-r. So asking for a defense of not-r would amount to committing straw man. But this might not be what is expressed by Peter’s “Why not?” Peter might only ask for a justfication of Olga’s retraction of commitment to the proposition that the good skating weather is holding. In that case the profile would be better depicted as follows:

- P: p

- O: Why(p)?

- P: r

- O: Why(r)?

- P: Why(Why(r)?)?

4 Many Questions

Though there are no doubt earlier uses of the idea of profiles of dialogue to be found, to my knowledge the first use of them going by that name occurs in Douglas Walton’s discussion of the spouse beating question (Walton 1989: 68,69, 1989a: 37,38). What is wrong with the question: “Have you stopped beating your spouse?” Traditionally, this question is supposed to exemplify a fallacy: the Fallacy of Many Question; for it seems that both, an affirmative answer and a denial, imply a confession of having beaten one’s spouse. But Walton shows that nothing need to be objectionable about this question if it occurs at the appropriate place in a normative profile of dialogue. Some questions ought be asked first to establish that the presuppositions of the spouse beating question are among the commitments of the answerer: “Do you have a spouse?”, “Have you ever beaten your spouse?”, and perhaps: “Did you do so regularly?”. If the answers to all these questions are in the affirmative, the questioner is entitled to asking the spouse beating question.

Questions and answers are presented by Walton in a (concrete) tree-shaped profile; each branch of the profile displaying an acceptable (nonfallacious) dialogue. The fallacious uses of the spouse beating question can then be explained as cases in which the questioner skips parts of a branch in the profile and thus creates a new branch that was not displayed as a legal possibility.

Profiles of dialogues have been used to study a number of other fallacies besides many questions: ignoratio elenchi (Krabbe 1992); the argumentum ad ignorantiam (Krabbe 1995, Walton 1996, 1999); non sequitur (Krabbe 1995a, 1996). In this paper I want to put them to use for the study of equivocation. First, however, I shall show how profiles can be used to depict the basic structure of the initial part of a critical discussion.

5 As discussion starts

The following dialogue displays a legitimate alternative to Peter’s “Why not?” move:

- Peter: The fine skating weather is holding.

- Olga: Why?

- Peter: Well, maybe not.

In an abstract profile:

- P: p

- O: Why(p)?

- P: No commitment(p)

Even though no argument for p has been put forward, and the dialogue is very brief, the purpose of critical discussion has been attained. Once he became aware of Olga’s critical position, Peter realized that by lack of arguments he had better throw in the towel. So he changed his position, and the original conflict was thereby resolved.

In the next example, Peter actually puts forward an argument. This time it is Olga who gives in.

- Peter: The fine skating weather is holding.

- Olga: Why?

- Peter: The almanac says so.

- Olga: I see.

This yields the following abstract profile:

- P: p

- O: Why(p)?

- P: r

- O: Concession(p)

But of course, Olga may be more critical. She may think that there is no good reason to believe what is in the almanac:

- Peter: The fine skating weather is holding.

- Olga: Why?

- Peter: The almanac says so.

- Olga: So what?

By “So what?” Olga expresses connection criticism of Peter’s argument: she critically questions whether the reason stated in Peter’s argument (“The almanac says so”) connects in a adequate and supportive way to the conclusion of that argument (“The fine skating weather is holding”). In a abstract profile:

- P: p

- O: Why(p)?

- P: r

- O: Why(r⇒p)?

It may also be that Olga, far from being skeptical about the almanac, would, on the contrary, be eager to accept Peter’s argument, if only its premise were true. She had a quick glance at the almanac herself, but does not remember what it said about the weather for the days to come.

- Peter: The fine skating weather is holding.

- Olga: Why?

- Peter: The almanac says so.

- Olga: Where?

By “Where?” Olga expresses tenability criticism. She now questions the tenability of the premise itself, apart from its connection with the conclusion of Peter’s argument. In the abstract profile, the question “Where?” reduces to a “Why?”:

- P: p

- O: Why(p)?

- P: r

- O: Why(r)?

Tenability and connection criticism correspond exactly to Næss’s dichotomy of tenability and relevance (Næss 1966, Krabbe 1987). They are the two main types of contesting an argument such that the oppponent (the critic) need not take upon herself any burden of proof. Two other types of criticism are active criticism and fallacy criticism. In both cases the opponent claims that there is something wrong with the argument, but only in the latter case is it claimed that the argument is inadmissible and should, therrefore, be retracted. In active criticism the proponent’s argument may be admitted as expressing a legitimate point of view, but then the opponent goes on to show why it is not sufficient for acceptance of the conclusion. For instance, the opponent may give a counterargument; that is, an argument for the opposite conclusion. (Clearly, when the opponent argues for a point of view, she has become a proponent of that point of view; that is, a reversal of roles has taken place. It is time to start using “W” and “B” instead of “P” and “O”.) The following example displays a case of active criticism, in fact of counterargument:

- Peter: The fine skating weather is holding.

- Olga: Why?

- Peter: The almanac says so.

- Olga: But the radio announced a storm.

Its abstract profile:

- W: p

- B: Why(p)?

- W: r

- B: s; therefore not-p

Here B’s last move contains an argumentative text, which can again be analyzed dialectically. It contains the following implicit dialogue:

- B: not-p

- W: Why(not-p)?

- B: s

An example of fallacy criticism will be presented in the next section.

6 Profiles for equivocation criticism

According to the immanent dialectical view on fallacies, the status of a move or argument must be decided in discussion, by the participants themselves (Krabbe, 1999). In fallacy criticism it is upon the critic to show why an alleged move in critical discussion is so completely wrong that it cannot even prima facie be accepted as a serious contribution to the discussion. Thus fallacy criticism leads to a critical discussion on a second level, a discussion about the permissibility of a move in the ground level discussion. Here the specter of a hierarchy of levels looms heavily. Though one may take comfort from the observation that in practice the disputants will never move up more than a few levels, many-leveled dialogue systems, such that (alleged) fallacies on one level are discussed on the next, are indeed quite complicated (Mackenzie 1979, 1981).

The method of profiles can also be useful as a heuristics for the development of such systems. A first step is to gain insight in the structure of discussions about various types of fallacies. Let us, by way of example, take a look at a case of a charge of equivocation:

- Peter: The fine skating weather is holding.

- Olga: Why?

- Peter: The weather will be good for sports.

- Olga: That’s no way to argue!

- Peter: What’s wrong about it?

- Olga: Well, there is this ambiguity.

- Peter: Which one?

- Olga: It depends on what you mean by “sports”. Do you mean “all sports” or “some sports”? In the first case, you should not say that the weather will be good for all sports, since obviously that cannot be the case. In the second case, your statement that the weather will be good for some sports will clearly not lead to the conclusion that it will be good for skating.

- Peter: By “sports” I mean “winter sports”.

The corresponding abstract profile of dialogue runs as follows:

- W: p

- B: Why(p)?

- W: r

- B: Inadmissible!

- W: Why(Inadmissible!)?

- B: Equivocatio!

- W: Why(Equivocatio!)?

- B: T=T_{1} or T=T_{2}; if T=T_{1} then r inadmissible; if T=T_{2} then (r⇒p) inadmissible - W: T=T_{3}

Confronted with the argument “r; therefore p”, B reacts with a charge of fallacy (“Inadmissible!”). This leads to a role reversal. In the ensuing second order discussion B (the original opponent) acts as the proponent of the fallacy charge, whereas W (the original proponent) acts as its opponent (“Why(Inadmissible!)?”). B then specifies what fallacy has been committed (Equivocatio!), and W challenges that specification (Why(Equivocatio!)?). In the next move B substantiates the charge. Next, there are many ways in which W can defend herself: she could opt for T=T_{1} and challenge the alleged inadmissibility of r under that interpretation, or for T=T_{2} and challenge the alleged inadmissibility of (r⇒p). She also could challenge the qualification of “equivocation” on the ground that T_{1} (“all sports”) and T_{2} (“some sports”) are not bona fide terms. Consequently, so she might plead, if there is a fallacy of ambiguity here, it would have to be amphiboly, not equivocation. Thus she could file a countercharge of fallacy and enter a discussion on the third level. In the present case, however, she offers a third interpretation of the allegedly equivocal term (T=“sports”), but her precization of r (r′=“The weather will be good for winter sports”) is unlikely to get her out of trouble.

Another option for W is to give in, and to retract reason r. She may then either retract her commitment to p altogether or go on to defend p in some other way. One subtle issue here is whether supplanting r by a precization such as r′ would count as a retraction. The question is whether in that case W admits that a fallacy has been committed (and is willing to pay the fine of $1,- that goes with committing a fallacy). If so, the precization would count as a retraction, but W might also maintain that her precization r′ was a reasonable way to understand r right from the start, in which case she still may plead not guilty.

These various moves and countermoves can all be displayed in a tree-shaped normative profile of dialogue (see Figure 1). As a last example I propose the following somewhat subtler case of equivocation:

- Peter: The fine skating weather is holding.

- Olga: Why?

- Peter: The almanac is right.

- Olga: Why?

- Peter: We are having fine skating weather right now.

- Olga: That’s no way to argue!

- Peter: What’s wrong about it?

- Olga: Well, there is this ambiguity.

- Peter: Which one?

- Olga: It depends on what you mean by “being right”. Does it mean “being sometimes right” or “being always right”? The present fine skating weather only shows that the almanac is sometimes right. But in order to show that the fine skating weather is holding you must interpret “being right” as “being always right”, or at least as “being right about tomorrow”. For, even if the almanac happens to be right about today, it may be wrong about tomorrow!

This is a subtler case of equivocation, because, as the abstract profile makes clear, two arguments are involved.

- W: p

- B: Why(p)?

- W: r

- B: Why(r)?

- W: s

- B: Inadmissible!

- W: Why(Inadmissible!)?

- B: Equivocatio!

- W: Why(Equivocatio!)?

- B: T=T_{1} or T=T_{2}; if T=T_{1} then the argument “r; therefore p” is inadmissible; if T=T_{2} then the argument “s, therefore r” is inadmissible

Here T=being right, T_{1}=being sometimes right, and T_{2}=being always right.

W: p

B: Why(p)?

* W: r

B: Why(r)? (tenability criticism)

B: Why(r⇒p)? (connection criticism)

B: s; there- fore not-p (active criticism)

B: Inadmissible! (fallacy criticism & role reversal)

W: Why(Inadmissible!)?

B: Equivocatio!

W: Why(Equivocatio!)?

B: T=T_{1} or T=T_{2}; if T=T_{1} then r inadmissible; if T=T_{2} then (r⇒p) inadmissible

W: T=T_{1}; Why(r inadmissible)?

W: T=T_{2}; Why((r⇒p)inadmissible)?

W: T=T_{3} W: Retraction (move *) etc.

B: Ignoratio elenchi! B: Non sequitur! etc.

B: Concession(p)

W: No commitment(p)

Figure 1: A normative profile of dialogue

7 Conclusion

I hope the foregoing examples have given some plausibility to the thesis that profiles of dialogues can be used as instruments for the analysis and the evaluation of discussions. They seem especially useful where discussions become complex and display such features as role reversal, active criticism, or fallacy criticism. Though this paper was concerned mainly with explicit discussions, the method of profiles might also be of use for the study of argumentative texts in which discussions remain implicit. In fact, one implicit discussion that occurred within an explicit discussion was analyzed by the profile method at the end of Section 5. Finally, the method of drawing up various profiles may serve as a heuristics for the development of fully formalized dialogue models (dialectic systems) that will account for more features of discussion than do extant systems.

Acknowledgement

I gratefully acknowledge to have profited from technical assistance by Hauke de Vries and Barteld Kooi in preparing this text and Figure 1.

References

Barth, Else M. & Erik C.W. Krabbe, 1982: From Axiom to Dialogue: A Philosophical Study of Logics and Argumentation. Berlin & New York: Walter de Gruyter.

Krabbe, Erik C.W., 1987: ‘Næss’s Dichotomy of Tenability and Relevance.’ In: Frans H. van Eemeren, Rob Grootendorst, J. Anthony Blair and Charles A. Willard (eds.), Argumentation: Across the Lines of Discipline: Proceedings of the Conference on Argumentation 1986, Dordrecht and Providence, RI: Foris Publications, pp. 307-16.

Krabbe, Erik C.W., 1992: ‘So What? Profiles for Relevance Criticism in Per-

suasion Dialogues,’ Argumentation 6 (2), 271-83. Krabbe, Erik C.W., 1995: ‘Appeal to Ignorance.’ In: Hans V. Hansen &

Robert C. Pinto (eds.), Fallacies: Classical and Contemporary Readings. University Park, PA: The Pennsylvania State University Press, pp. 274-86.

Krabbe, Erik C.W., 1995a: ‘Can We Ever Pin One Down to a Formal Fallacy?’ In: Frans H. van Eemeren, Rob Grootendorst, J. Anthony Blair and Charles A. Willard (eds.), Proceedings of the Third ISSA Conference on Argumentation (University of Amsterdam, June 21-24, 1994) II: Analysis and Evaluation, Amsterdam: Sic Sat, International Centre for the Study of Argumentation, pp. 333-344. Also in: Theo A. F. Kuipers and Anne Ruth Mackor (eds.), Cognitive Patterns in Science and Common Sense: Groningen Studies in Philosophy of Science, Logic, and Epistemology, Amsterdam and Atlanta, GA: Rodopi, 1995, pp. 151-64.

Krabbe, Erik C.W., 1996: ‘Can We Ever Pin One Down to a Formal Fallacy?’ In: Johan van Benthem, Frans H. van Eemeren, Rob Grootendorst and Frank Veltman (eds.), Logic and Argumentation, Amsterdam, etc.: North-Holland, 1996, pp. 129-141 (reprint of Krabbe 1995).

Krabbe, Erik C.W., 1999: ‘The Dialectic of Quasi-Logical Argument.’ To be published in the proceedings of the Fourth ISSA-Conference on Argumentation (Amsterdam, 16-19 June 1998).

Mackenzie, Jim D. 1979: ‘How to Stop Talking to Tortoises,’ Notre Dame

Journal of Formal Logic 20, 705-17. Mackenzie, Jim D. 1981: ‘The Dialectics of Logic,’ Logique et analyse n.s. 24,

159-77. Næss, Arne, 1966: Communication and Argument: Elements of Applied Seman-

tics. Oslo: Universitetsforlaget, and London: Allen & Unwin. Translation of: En del elementære logiske emner, Oslo, 1947, etc.

Van Eemeren, Frans H. & Rob Grootendorst, 1992: Argumentation , Communication, and Fallacies: A Pragma-Dialectical Perspective. Hillsdale, NJ, & Hove & London: Lawrence Erlbaum.

Walton, Douglas N., 1989: Question-Reply Argumentation. Greenwood Press:

New York. Walton, Douglas N., 1989a: Informal Logic: A Handbook for Critical Argumen-

tation. Cambridge U.P.: Cambridge. Walton, Douglas N., 1996: Arguments from Ignorance. University Park, PA:

The Pennsylvania State University Press. Walton, Douglas N., 1999: ‘Profiles of Dialogue for Evaluating Arguments from

Ignorance,’ Argumentation 13 (1), 53-71. Walton, Douglas N. & Erik C.W. Krabbe, 1995: Commitment in Dialogue:

Basic Concepts of Interpersonal Reasoning. Albany, NY: State University of New York Press.